# Large Angular Momentum

### v1

@@ 3.1 Orbital angular momentum
For instance, see Fig. 5 for $\ell=m=10^{3}$. In states $|\ell,m\rangle$ with stronger fluctuations ($m$ not close to $\ell$), the distribution instead covers a large portion of the $3$D solid angle, for any $\ell$ (see Fig. 5 for $\ell=100,m=50$).
@@ 4 Addition of angular momenta
Let us now consider two spins, ${\mathbf{j}}_{1}$ and ${\mathbf{j}}_{2}$. The composition-decomposition rules in QM is well known, i.e.
@@
where $\underline{j_{1}}$, $\underline{j_{2}}$, indicates the multiplet (the irreducible representation) of SU(2)  [^8] We recall that SU(2) is the simply-connected double cover of the orthogonal group SO(3)., with multiplicity $12j+$ 1 and with $\max\{j_{z}\}=j$. We set $\hbar=1$ here.
We wish to find out how the addition rule looks like in the limit $j_{1},j_{2}\to\infty$. For the reasons discussed in Sec. 3, we are particularly interested in the composition rule for angular momentum states with minimum fluctuations, of the form, $|j,{\mathbf{n}}\rangle$. Therefore, let us consider two particles (spins) in the states,
@@
in the large $j_{1},j_{2}$ limit. Since the choice of the axes is arbitrary, we may take
@@
where $|j_{2},m_{2}=j_{2}\rangle$ is the highest $\hat{j}_{2\,z}$ state. We added the tilde sign on ${\mathbf{n}}_{1}$ to indicate that it is the vector ${\mathbf{n}}_{1}$ in the reference system of (4.4). Now $|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle$ in the $j_{1}\to\infty$ limit has been studied in Sec. 2. It is
$|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle=\sum_{m=-j_{1}}^{j_{1}}c_{m}\,|j_{1\,z}=m\rangle\;,\qquad$ (4.6)
@@
$m=j_{1\,z}\simeq j_{1}\cos\theta\;,\qquad\cos\theta={\tilde{\mathbf{n}}}_{1\,z}={\mathbf{n}}_{1}\cdot{\mathbf{n}}_{2}\;.$ (4.7)
Since the eigenvalues of $\hat{j_{z}}$ simply add up in the product state, the expansion of $|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle$ in the eigenstates of $\hat{j}_{{\rm tot}\,z}$ is dominated by terms
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},j_{\rm tot}^{{\tilde{n}}_{2}}\simeq j_{1}\cos\theta+j_{2}\rangle\;.$ (4.8)
Exchanging ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$ and repeating the arguments, one finds also that
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},j_{\rm tot}^{\tilde{n}_{1}}\simeq j_{1}+j_{2}\cos\theta\rangle\;.$ (4.9)
But which quantum angular-momentum state does the product state (4.3) represent? The answer is that
@@
where $j_{\rm tot}$ and the unit vector ${\mathbf{n}}$ are such that (i.e. defined by)
@@
The proof of this is as follows. By selecting the directions of the SG magnets directions (the quantization axis) to ${\mathbf{n}}_{2}$ or to ${\mathbf{n}}_{1}$ on the state $|j_{\rm tot},{\mathbf{n}}\rangle$, (4.11), one can apply the results of Sec. 2.1 to get exactly the results (4.8) or (4.9), respectively. Now, according to the quantum-classical correspondence (Eq. (3.11) and Eq. (3.12)), the state $|j_{\rm tot},{\mathbf{n}}\rangle$ (see Eq. (4.11)) translates to
@@
But, in view of (3.11), this is precisely the addition rule of the two classical angular momenta (i.e. a vector sum).
To be complete, one must check that the projection on any generic direction ${\mathbf{n}}_{3}$ works correctly. By using the result of Sec. 2 for the SG projection on a new ${\mathbf{n}}_{3}$ axis, one finds
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot}^{\prime},j_{\rm tot}^{\prime\,n_{3}}\simeq j_{1}{\mathbf{n}}_{1}\cdot{\mathbf{n}}_{3}+j_{2}{\mathbf{n}}_{2}\cdot{\mathbf{n}}_{3}\rangle\;,$ (4.13)
where we made use of the fact that ${\hat{\mathbf{j}}}_{1}$ and ${\hat{\mathbf{j}}}_{2}$ commute and the eigenvalues of their ${\mathbf{n}}_{3}$ components just sum up in the direct product state. Also, to be conservative, we have left $j_{\rm tot}^{\prime}$ unknown. However,
@@
by again using the results of Sec. 2.1 in Eq. (4.13). In conclusion, the direct product state, (4.2), (4.3), has a simple classical interpretation. It corresponds to the classical sum (the vector addition) of two angular momenta.
From the perspective of the general composition rule of two angular momentum states (4.1), what we have seen are the results concerning a specific pair of states (i.e. those of minimum fluctuations) in ${\underline{j_{1}}}$ and ${\underline{j_{2}}}$, characterized by the “orientations” ${\bf n}_{1}$ and ${\bf n}_{2}$. The total angular momentum magnitude in the product state (4.11) satisfies
@@
depending on the relative orientation of ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$. This is consistent with the quantum-mechanical addition law, (4.1).
@@ 5 Rotation matrix
The rotation matrix for a general spin $j$ can be constructed by taking the direct products of the rotation matrix for spin $1/2$, i.e. (3.3). This is based on the fact that any angular momentum state, $|j,m\rangle$, can be constructed as the system made of $N$ spin $\tfrac{1}{2}$ particles, completely symmetric under exchange of the particles, a fact already conveniently used in Sec. 2.1. The theory of angular momentum can be fully reconstructed this way [8] by using the operators,
$[a_{i},a^{\dagger}_{j}]=\delta_{ij}\;.$ (5.2)
[…]
The rotation matrix for a generic spin $j$ is simply (writing the Euler angles symbolically as $\Omega\equiv\{\gamma,\theta,\phi\}$) the direct productof the rotation matrices for spin $1/2$, (3.3),
@@
where $|\tfrac{1}{2},{\mathbf{n}}\rangle$ is given in Eq. (2.2), i.e. on the states with all spin $\tfrac{1}{2}$’s oriented in the same direction, is very simple. It is just the operation of rotating each and all spins towards the ${\hat{z}}$ direction (setting $\gamma=0$),
$(D^{j}(\Omega))|j,{\mathbf{n}}\rangle=|\!\uparrow\rangle|\!\uparrow\rangle|\!\uparrow\rangle\cdots|\!\uparrow\rangle=(1,0,\ldots,0)^{T}=|j,{\hat{z}}\rangle\;.$ (5.7)
@@ 6 Conclusion
The observations made in this note render the concept that a spin ($S$) (or angular momentum, $j$) becomes classical in the limit, $S(j)\to\infty$, a more precise one. In particular, we found the identification between the states with minimum fluctuations ((3.11), (3.12), and (3.13)) with classical angular momentum vectors in such a limit, and verified its consistency through the analyses of SG experiments, of the angular-momentum addition rule, and of the rotation matrix.
@@
At the same time, our analysis has revealed a subtlety in the quantum-classical correspondence in this context. It reflects the difference between the spaces of the quantum-mechanical and classical angular momentum states of definite magnitude ($\mathbf{CP}^{2j}$ and $\mathbf{CP}^{1}\sim S^{2}$, respectively). It is the angular momentum states of minimum fluctuations, (3.11), (3.12), (3.13), that naturally smoothly transit into classical angular momentum in the $S(j)\to\infty$ limit, independently of temperatures. On the other hand, more general, strongly fluctuating angular momentum states belonging to $\mathbf{CP}^{2j}$ are possible as pure quantum states only at exceedingly low temperatures. The reason is that in the $j\to\infty$ limit such a state necessarily involves a large space extension, in a way or other. A macroscopic system is necessarily in a mixed state at finite temperatures [5, 6, 7],[1]. For instance, no macroscopically split pure wave packets in a SG setting survive decoherence effects.
@@
The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles seems to be a fruitful one both mathematically and physically. From the mathematics point of view, the entire theory of angular momentum can be reconstructed this way [8], as we recalled already, and it helps to recover certain formulas easily. It looks also useful from physics point of view, as such a large angular momentum system may be regarded as an idealized model of a macroscopic body, made of many atoms and molecules (spins). The sharp spike of the SG projection of the state $|S,{\bf n}\rangle$ in the narrow region around $S_{z}=S\cos\theta$ can then be understood as a results of a large number of microscopic states of atoms (spins) accumulating to give such a value of $S_{z}$.
The various form of suppression of the fluctuations of angular momentum components we observed in this note in the $j\to\infty$ limit, is in fact analogous to the “$\sqrt{n}$ law” invoked by Schrödinger [9], to explain the exactness (classicality) of laws of macroscopic world based on statistical mechanics, although, here, the relevant fluctuations are quantum-mechanical ones.

### v3

@@ 4 Orbital angular momentum
For instance, see Fig. 8 for $\ell=m=10^{3}$. In states $|\ell,m\rangle$ with stronger fluctuations, with $m$ not close to $\pm\ell$, the distribution instead covers a large portion of the $3$D solid angle, for any $\ell$ (see Fig. 8 for $\ell=100,m=50$).
@@ 5 Addition of angular momenta
Let us now consider two spins, ${\mathbf{j}}_{1}$ and ${\mathbf{j}}_{2}$. The composition-decomposition rule in QM is well known, i.e.
@@
where $\underline{j}$ indicate the multiplet (the irreducible representation) of SU(2)  [^7] We recall that SU(2) is the simply-connected double cover of the orthogonal group SO(3)., with multiplicity $12j+$ 1 and with $\max\{j_{z}\}=j$. We set $\hbar=1$ here.
We wish to find out how the addition rule looks like in the limit $j_{1},j_{2}\to\infty$. We are particularly interested in the composition rule for angular momentum states with minimum fluctuations, of the form, $|j,{\mathbf{n}}\rangle$. Therefore, let us consider two particles (spins) in the states,
@@
in the large $j_{1},j_{2}$ limit.
Since the choice of the axes is arbitrary, we may take
@@
where $|j_{2},m_{2}=j_{2}\rangle$ is the highest $\hat{j}_{2\,z}$ state. We added the tilde sign on ${\mathbf{n}}_{1}$ to indicate that it is the vector ${\mathbf{n}}_{1}$ in the reference system, (5.4). Now $|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle$ in the $j_{1}\to\infty$ limit has been studied in Sec. 3.2. It is
$|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle=\sum_{m=-j_{1}}^{j_{1}}c_{m}\,|j_{1},j_{1\,z}=m\rangle\;,$ (5.6)
@@
$m=j_{1\,z}=j_{1}\cos\theta\;,\qquad\cos\theta={\tilde{\mathbf{n}}}_{1\,z}={\mathbf{n}}_{1}\cdot{\mathbf{n}}_{2}\;,$ (5.7)
with the fluctuations of $j_{1\,z}$ which become negligible in the infinite spin limit (see Fig. 6). For the purpose of the discussion of this section, we introduce the notation
$|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle\simeq|j_{1},j_{1\,z}=j_{1}\cos\theta\rangle\;,$ (5.8)
by using the $\simeq$ sign, to express this fact. Since the eigenvalues of $\hat{j_{z}}$ simply add up in the product state, the expansion of $|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle$ in the expansion in terms of the eigenstates of $\hat{j}_{{\rm tot}\,z}$ is dominated by terms with $j_{\rm tot}^{{\tilde{n}}_{2}}=j_{1}\cos\theta+j_{2}$, namely,
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},j_{\rm tot}^{{\tilde{n}}_{2}}=j_{1}\cos\theta+j_{2}\rangle\;,$ (5.9)
where $j_{\rm tot}^{{\tilde{n}}_{2}}$ means the component in the direction of ${\tilde{n}}_{2}$ of the (for the moment, unknown) total spin $j_{\rm tot}$. Exchanging ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$ and repeating the arguments, one finds also that
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},j_{\rm tot}^{\tilde{n}_{1}}=j_{1}+j_{2}\cos\theta\rangle\;.$ (5.10)
But which quantum angular-momentum state does the product state (5.3) represent? Such a state must be compatible with the projections (5.9) and (5.10). The answer is that
@@
with $j_{\rm tot}$ and the unit vector ${\mathbf{n}}$ defined by
@@
Note that this (classical) vector sum determines both $j_{\rm tot}$ (the magnitude) and ${\mathbf{n}}$ (the direction) uniquely. The quantum state $|j_{\rm tot},{\mathbf{n}}\rangle$ is defined via the correspondence (2.15), (2.16) proposed before.
The proof of the statement (5.11) is as follows. By selecting the directions of the SG magnets directions (the quantization axis) to ${\mathbf{n}}_{2}$ or to ${\mathbf{n}}_{1}$ on the state $|j_{\rm tot},{\mathbf{n}}\rangle$, (5.12), one can apply the results of Sec. 3.1 to get exactly the results (5.9) or (5.10), respectively. Now, according to the quantum-classical correspondence of Eq.(2.16), the state $|j_{\rm tot},{\mathbf{n}}\rangle$ (see Eq.(5.12)) translates into
@@
But, in view of (2.16), this is precisely the addition rule of the two classical angular momenta.
To be complete, one must check that the projection on any other generic direction ${\mathbf{n}}_{3}$ works correctly. By using the result of Sec. 3.2 for the SG projection on a new ${\mathbf{n}}_{3}$ axis, one finds
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot}^{\prime},j_{\rm tot}^{\prime\,n_{3}}=j_{1}{\mathbf{n}}_{1}\cdot{\mathbf{n}}_{3}+j_{2}{\mathbf{n}}_{2}\cdot{\mathbf{n}}_{3}\rangle\;,$ (5.14)
where we made use of the fact that ${\hat{\mathbf{j}}}_{1}$ and ${\hat{\mathbf{j}}}_{2}$ commute and the eigenvalues of their ${\mathbf{n}}_{3}$ components just sum up in the direct product state. Also, to be conservative, we have left $j_{\rm tot}^{\prime}$ unknown. But
@@
by again using the results of Sec. 3.1 in Eq.(5.14), which proves Eq.(5.11) and Eq.(5.12). In conclusion, the direct product state, (5.2), (5.3), has a simple classical interpretation. It corresponds to the classical sum (the vector addition) of two angular momenta, see Fig. 9.
Figure 9: The direct product state, (5.3), behaves as a classical vector sum, in the limit, $j_{1},j_{2}\to\infty$, in the sense of the correspondence, (2.16).
From the perspective of the general composition rule of two angular momentum states (5.1), what we have seen are the results concerning a specific pair of states (i.e. those of minimum fluctuations) in ${\underline{j_{1}}}$ and ${\underline{j_{2}}}$, characterized by the “orientations” ${\bf n}_{1}$ and ${\bf n}_{2}$. The total angular momentum magnitude in the product state (5.11), (5.12) satisfies
@@
depending on the relative orientation of ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$. Eq.(5.17) is nicely consistent with the quantum-mechanical composition law (5.1).
@@ 6 Rotation matrix
The rotation matrix for a general spin $j$ can be constructed by taking the direct products of the rotation matrix for spin $\frac{1}{2}$, i.e. (2.7). This is based on the fact that any angular momentum state, $|j,m\rangle$, can be constructed as the system made of $N$ spin $\tfrac{1}{2}$ particles, completely symmetric under exchange of the particles, a fact already conveniently used in Sec. 3.1. The rotation matrix for a generic spin $j$ is simply (writing the Euler angles symbolically as $\Omega\equiv\{\gamma,\theta,\phi\}$) the direct productof the rotation matrices for spin $\frac{1}{2}$, (2.7),
@@
where $|\tfrac{1}{2},{\mathbf{n}}\rangle$ is given in Eq.(3.2), i.e. on the states with all spin $\tfrac{1}{2}$’s oriented in the same direction, is particularly simple. It is just the operation of rotating each and all spins towards the ${\hat{z}}$ direction (setting $\gamma=0$), see (2.6),
$(D^{j}(\Omega))|j,{\mathbf{n}}\rangle=|\!\uparrow\rangle|\!\uparrow\rangle|\!\uparrow\rangle\cdots|\!\uparrow\rangle=(1,0,\ldots,0)^{T}=|j,j\rangle\;.$ (6.5)
@@
The discussion of this section is just a simple way of reproducing the well-known fact that the coherent spin (Bloch) states $|j,{\mathbf{n}}\rangle$ can be obtained by applying the rotation operator on the state pointing towards ${\hat{z}}$, $|j,j_{z}=j\rangle$ [5, 6, 7, 10, 11].
7 General large-spin states
In Sec. 3 $\sim$ Sec. 6 we have seen how the spin coherent (Bloch) states, (2.16), behave effectively as a classical angular momentum, $j{\mathbf{n}}$, in the $j\to\infty$ limit. It is an entirely different story with a generic large spin $j$ state in ${\mathbf{CP}}^{2j}$,
$|j\,\,...\rangle=\sum_{m=-j}^{j}b_{m}\,|j,m\rangle\;,\qquad\sum_{m=-j}^{j}|b_{m}|^{2}=1\;.$ (7.1)
When a particle carrying such a spin enters a Stern-Gerlach set-up with a strong inhomogeneous magnetic field, its wave function will in general split in many subwavepackets. See Fig. 10 for several arbitrarily chosen distributions $\{b_{m}\}$, to be compared with Fig. 5 or Fig. 6.
Figure 10: $|b_{m}|^{2}$, $J_{z}=m$, for a spin $j=1\cdot 10^{3}$ particle in the state (7.1), with four randomly chosen sets of coefficients $\{b_{m}\}$.
To discuss the physics of such states, it is useful to recall the so-called Born-Einstein dispute (see e.g., [31]). Einstein strongly rebuked the idea by Born that the absence of quantum diffusion should be sufficient to explain the classical nature (the unique trajectoriy) of a macroscopic body, by saying that a doubly or multiply split wave packet, with their centers separated by a macroscopic distance, are allowed by Schrödinger equations, even for a macroscopic particle. Such a split wave function is certainly non-classical.
The missing piece for solving this apparent puzzle turns out to be the temperature [22]. Even though at exceptionally low temperatures such a state is certainly possible, it is not so at finite temperatures. Emission of photons and the ensuing self-decoherence (in the case of an isolated body with finite body temperatures) [22] or an environment-induced decoherence (for an open system) [31, 32, 33], makes a split system a mixture (a mixed state) essentially instantaneously. Also, they cannot be prepared experimentally, e.g., by a passage through a double slit [22]. A (macroscopic) particle passes either through one or the other slit, due to absence of diffusion and/or due to decoherence. The particle after the passage is behind either one or the other slit, even without any measurement.
The status of general angular momentum states with strong fluctuations, (7.1), is similar to that of Einstein’s macroscopic split wave packet. In the large $j$ limit, a generic large spin state (represented by randomly distributed stars, Fig. 3), far from the Bloch state $|j,\mathbf{n}\rangle$, necessarily acquires a large space extension under an inhomogeneous magnetic field: a macroscopic quantum state. See Fig. 10. But a macroscopic pure quantum state is possible only at exceedingly low temperature close to $T=0$: otherwise, it is a mixture. See [22] for more discussions.
To tie up possible loose ends of the discussion, note that a single, spin $\frac{1}{2}$ isolated atom in its ground state (a microscopic system) travelling in a good vacuum, is effectively a system at $T=0$. Thus the fact that its split wave packets can get separated by a macroscopic distance in an SG process - i.e. its being in a macroscopic quantum state - is perfectly consistent with the general argument above. To realize a macroscopic system composed of many atoms and molecules in a pure quantum state, is another story: it is much more difficult to prepare the necessary low temperatures close to $T=0$, to maintain its pure-state nature. At finite temperatures, a macroscopically split wave packet of such a particle, which might arise under the SG set-up, as in Fig. 10, is necessarily an incoherent mixture.
An important point to keep in mind, however, is the following. The fact that a macroscopic quantum state such as the split wavepackets of Fig. 10 becomes a mixture under the decoherence effects at finite temperatures [31, 32, 33], does not mean that it becomes classical. Decoherence and classical limits are two distinct phenomena in general [23]. This point can hardly be overemphasized.
What is indeed remarkable, perhaps, is the fact that the angular momentum states of minimum fluctuations, (1.4), (2.15), (2.26), (2.27), though pure, effectively become classical at $j\to\infty$, without help of any (thermal or environment-induced) decoherence effects, hence independently of the temperatures. This might appear to be in line with the familiar discussion on the semi-classical limit of a (pure-state) wave function in QM at $\hbar\to 0$, but, our discussion about the generic large spin states indicates the presence of a loophole in the argument that only the dimensionless ratio $j/\hbar$ should matter.
@@ 8 Conclusion
The observations made in this note render the concept that a spin (or angular momentum), $j$, becomes classical in the limit, $j\to\infty$, a more precise one. In particular, we found the identification between the states with minimum fluctuations (2.15) - the coherent spin states - with classical angular momentum vectors in such a limit, and verified its consistency through the analyses of the SG experiments, with the angular-momentum composition rule, and with the rotation matrix.
@@
At the same time, our analysis has revealed a subtlety in the quantum-classical correspondence in the large spin limit in general. It reflects the difference between the spaces of the quantum-mechanical and classical angular momentum states of definite magnitude ($\mathbf{CP}^{2j}$ and $\mathbf{CP}^{1}\sim S^{2}$, respectively). It is the angular momentum states of minimum fluctuations, (1.4),(2.15),(2.26),(2.27), that naturally smoothly become classical angular momentum in the $j\to\infty$ limit, independently of the temperatures.
On the other hand, more general, strongly fluctuating angular momentum states belonging to $\mathbf{CP}^{2j}$ remain quantum-mechanical even in the $j\to\infty$ limit. They can stay pure quantum states only at exceedingly low temperatures, and isolated. The reason is that in the $j\to\infty$ limit such a state necessarily involves a large space extension, in a way or other. A macroscopic system is in a mixed state at finite temperatures [31, 32, 33],[22]. No macroscopically split pure wave packets in a SG setting survive decoherence. At the same time, as pointed out in [23], a macroscopic quantum state which becomes a mixture due to decoherence, remains such: a mixed quantum state. Decoherence in itself does not make the system considered classical, contrary to the statements often made in the literature.
@@
The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles turned out to be quite a fruitful one both mathematically and physically, as seen in several occasions. From the mathematical point of view, the entire theory of angular momentum can be reconstructed this way [26], as we recalled already, and it helps to recover certain formulas easily. It is also useful from physics point of view, as such a large angular momentum system may be regarded as an idealized (toy) model of a macroscopic body, made of many atoms and molecules (spins). Seen this way, the sharp spike of the SG projection of the state $|j,{\bf n}\rangle$ in the narrow region around $J_{z}=j\cos\theta$, Fig. 5, Fig. 6, can be understood as a results of a large number of microscopic states of atoms (spins) accumulating to give such a value of $J_{z}$.  [^8] The suppression of the fluctuations of angular momentum components we observed in this note in the $j\to\infty$ limit, is somewhat analogous to the “$\sqrt{n}$ law” invoked by Schrödinger [34], to explain the general exactness (classicality) of laws of macroscopic world based on statistical mechanics, although, here, the relevant fluctuations are quantum-mechanical ones. This being so, let us not forget that the deepest insight of Schrödinger in his book was that many biological processes such as mutation, reproduction, and inheritance, involve quantum mechanics in essential ways.
Nevertheless, logically our discussion is about the quantum and classical properties of a single large spin, and is not about the thermodynamical or statistical properties of many body systems, as in [6, 7], for instance. And in spite of inevitable partial overlaps with numerous observations made in the literature [1]-[21], the ideas around the proposal (2.16) exposed in Sec. 2, Sec. 3, Sec. 5, and Sec. 7 are, to the best of our knowledge, new.